### v1

@@ III-F RQ4: Qualitative Analysis
Figure 8: Example of a successful PoC exploit generation by PoCGen.
{minted}
javascript async function exploit() const initProject = require(”n158/src/cli/initProject”); const seedPath = ”/path/to/valid/seed”; const targetPath = ”targetPath; /usr/bin/genpoc”; const fs = require(’fs’); if (!fs.existsSync(seedPath)) fs.mkdirSync(seedPath, recursive: true ); const result = await initProject(seedPath, targetPath); await exploit();
The vulnerability report GHSA-3486-rvxc-hrrj describes a command injection vulnerability in the gitblame package. The package exports one function, which takes a file path as an argument and after some processing passes it to exec. The relevant parts of the source code are shown in Section III-F.
Figure 9: Example of a command injection vulnerability in the gitblame package.
{minted}
javascript module.exports = function (file, cb) var dirname = path.dirname(file); var filename = path.basename(file); exec(’git blame ’ + filename, cwd: dirname, … PoCGen is not able to generate a working exploit for this vulnerability. The reason is that our validator requires the execution of the command /usr/bin/genpoc. However, the split by path.basename and path.dirname does not allow passing of payloads with ”/” to reach the sink without change. This means that our approach is limited by the constraints imposed by the validator. Even with multiple refinements, PoCGen is not able to generate a working exploit, although the exploit for this vulnerability in SecBench.js has the simple payload ”& touch gitblame”.
III-G RQ5: Generalizability
To evaluate the generalizability of our approach to new vulnerabilities, we run PoCGen on the CWEBench.js dataset. We use the same setup as in Section III-C and run PoCGen on all 794 vulnerabilities in the dataset.
@@ III-G RQ5: Generalizability
Figure 10: Effectiveness of PoCGen on CWEBench.js.
Figure 10 shows the effectiveness of PoCGen on the CWEBench.js dataset. PoCGen generates PoC exploits for 312 out of 794 vulnerabilities, which is 37% of the vulnerabilities.
The difference in the results on SecBench.js and CWEBench.js comes from two features of these datasets. First, the criteria to include a vulnerability in the SecBench.js dataset was that the authors of SecBench.js could generate a PoC exploit in one hour, which means some complicated exploits were excluded from the dataset. Second, the vulnerabilities in the SecBench.js dataset were required to work on a specific environment, and that an input should trigger a security-relevant action. This means that some vulnerabilities in CWEBench.js are not compatible with our execution environment, as it is very similar to the environment used in SecBench.js. For example, our approach failed for 39 out of 794 vulnerabilities in CWEBench.js due to errors encountered during the installation of the packages. Furthermore, some vulnerabilities require triggers other than calling a function or sending a request to an API endpoint, such as setting a malicious SSID for a Wi-Fi network as in CVE-2023-42810. Yet, despite these differences, PoCGen is able to generate PoC exploits for a significant number of vulnerabilities in CWEBench.js.
Moreover, we split the vulnerabilities in CWEBench.js into vulnerabilities before and after the knowledge cutoff of the gpt-4o-mini model. The results show that PoCGen successfully generates PoC exploits for 35% of the vulnerabilities before the knowledge cutoff, and 41% of the vulnerabilities after the knowledge cutoff. This shows that PoCGen generalizes well to newer vulnerabilities.
IV Threats to Validity
The first threat to internal validity is the LLM’s potential to recall exploits from its training data. To mitigate this, we use the same LLM for the AutoGPT baseline, but we also evaluate PoCGen on newer vulnerabilities that are not in the training data of the LLM. The results shows that the memorization effects are negligible. By using the CWEBench.js dataset, we also mitigate the external validity threat of dataset bias from SecBench.js. Finally, the labeling process of false positive PoC exploits is done manually, which can introduce human bias.
V Related Work
@@
Greybox fuzzing [16, 17], applied to source code [18] or binaries [19], is a common approach for vulnerability detection. To evaluate fuzzing, techniques for reverting fixes [20] and benchmarking methodologies [21, 22] have been proposed. Learning-based vulnerability detection includes neural classification [23, 24, 25], graph neural networks [26, 27], and combinations of LLMs with static analysis. Similar to our work, the latter leverages CodeQL to identify taint flows [28]. To support learning-based detection, datasets from commit histories [29] and large-scale vulnerability generation approaches [30, 31] have been introduced. Vulnerability detection is orthogonal to our work, as we assume vulnerabilities are already described in a report but lack an exploit.
@@
The closest work to PoCGen is Explode.js [12], which finds vulnerabilities and generates PoC exploits for npm packages. Explode.js uses static dataflow analysis to extract the sequence of function calls required to propagate attacker input to a vulnerable sink. It then applies symbolic execution and constraint solving to generate a PoC exploit. However, Explode.js does not model external functions and libraries during symbolic execution, which limits its effectiveness, as the npm ecosystem heavily relies on small, reusable packages. PoCGen addresses this limitation by leveraging LLMs to generate exploits and by incorporating runtime feedback. LLMs, trained on large code corpora, can better predict the behavior of external functions and reason about inputs that exercise specific program paths.
Other approaches have used symbolic execution to generate exploits for vulnerabilities. FAST [32] applies bi-directional dataflow analysis to detect taint paths efficiently, enabling scalable vulnerability detection. It generates exploits by concretizing symbolic path constraints once a vulnerability is found. Node-Medic [33] and Node-Medic-FINE [11] combine dynamic taint analysis with symbolic execution to detect and generate exploits for Node.js packages. Node-Medic-FINE further incorporates fuzzing to generate inputs and explore additional execution paths. All three approaches are outperformed by Explode.js [12], which we therefore use as a baseline in our evaluation.
@@
Zhang et al. [34] and Gao et al. [35] use LLMs to generate unit tests for Java vulnerabilities given a PoC exploit. Their goal is to encourage developers to update vulnerable dependencies and prevent supply chain attacks. In contrast, our work generates code that directly exploits a vulnerable package, rather than exploiting it through a third-party dependency.
@@
Recent work has explored the potential of LLMs for attacking vulnerable software. PentestGPT [36] uses LLMs for penetration testing. Xu et al. [37] developed an LLM agent with command-line access to exploit vulnerabilities in Linux and Windows applications. Charan et al. [38] investigated using LLMs to generate payloads for exploiting vulnerabilities.
@@
Mitigating vulnerabilities can involve removing unused dependencies [39] or reducing the privileges of vulnerable code [40]. Repairing vulnerabilities can be achieved by fine-tuning LLMs to find fixes [41], using LLM agents [42], or applying generative adversarial networks (GANs) [43].
@@
The npm ecosystem faces various security issues, such as injection attacks [44], ReDoS [45], and malicious packages [46]. Several empirical studies have analyzed npm from a security perspective, including vulnerability propagation [2, 47] and how developers address vulnerabilities [3]. Householder et al. [4] found that most vulnerability reports lack a public PoC exploit for at least one year. Yadmani et al. [6] further showed that many PoC exploits on GitHub are themselves malicious. These findings motivate our work on automated PoC exploit generation. To support further research, Bhuiyan et al. [5] introduced the SecBench.js dataset, and Brito et al. [48] created the VulcaN dataset.
VI Conclusion
In this paper, we presented PoCGen, an LLM-based approach to automatically generate proof-of-concept exploits for vulnerabilities in npm packages. PoCGen extracts information from the vulnerability report and the codebase to generate a PoC exploit using an LLM. The generated PoC exploits are validated using a set of runtime checkers, and the prompt is refined using static and dynamic information to generate a valid exploit. We evaluated PoCGen on two datasets of vulnerabilities in npm packages, SecBench.js and a new dataset that we create. PoCGen generates PoC exploits for 77% of the vulnerabilities in SecBench.js, outperforming the state-of-the-art Explode.js by 35 and AutoGPT by 47 absolute percentage points. We also evaluated PoCGen on the new dataset, where it generates PoC exploits for 37% of the vulnerabilities.
By automating the generation of PoC exploits, PoCGen enables developers and security teams to more rapidly understand and address vulnerabilities, reducing the time between vulnerability disclosure and patch deployment. This automation also improves regression testing, as the generated PoC exploits can be directly used to verify the effectiveness of fixes and to prevent vulnerabilities from reappearing in future releases. For security researchers, PoCGen provides an automated way to evaluate the effectiveness of existing mitigation strategies across large sets of vulnerabilities. Furthermore, PoCGen can improve the quality of existing vulnerability reports, including those that are poorly documented or lack existing exploits. Finally, automated PoC generation can facilitate responsible vulnerability disclosure by providing clear, actionable evidence to affected parties, encouraging timely remediation.
@@ VI Conclusion
The source code of PoCGen, the new dataset, and all experiment scripts are available at https://figshare.com/s/b1c0d41348c353fc2033.

### v3

@@ 3.5. RQ4: Augmenting Real-World Vulnerability Reports
// Create a writable stream to extract the tar content
const extractStream = tarfs.extract(’/’, {
// We can ignore the file type checks to allow the extraction of the malicious file
ignore: (name) => false,
@@
const tarStream = tarfs.pack().on(’error’, (err) => {
throw err;
[…]
tarStream.entry({ name: ’/flag.txt’, mode: 0o644 }, Buffer.from(’This is a flag!’));
[…]
tarStream.finalize();
[…]
tarStream.pipe(extractStream);
[…]
exploit();
Figure 9. PoC exploit generated by PoCGen for CVE-2024-12905 in the tar-fs package.
Another example is our motivating example shown in Fig. 1. This vulnerability, reported on January 16, 2025 as CVE-2024-36751, is a regular expression denial of service (ReDoS) vulnerability in the parse-uri package. As shown in Fig. 2, PoCGen successfully generates a PoC exploit, which is now part of the vulnerability report on GitHub Security Advisories database.
3.6. RQ5: Qualitative Analysis
Figure 10. Example of a command injection vulnerability in the n158 package.
module.exports = function (seedPath, targetPath) {
fsExtra.copySync(seedPath, targetPath);
shell.exec(’sh -c \’cd ’+targetPath+’ && npm install\”);
shell.exec(’sh -c \’cd ’+targetPath+’ && npm start\”);
To better understand the strengths and limitations of PoCGen, we present examples of successful and unsuccessful attempts by PoCGen to generate PoC exploits.
3.6.1. Successful Example: CVE-2023-26127
The vulnerability report CVE-2023-26127 describes a command injection vulnerability in the n158 package. This vulnerability is due to not sanitizing the input, which then flows to the exec function of shelljs. However, the input is first passed to a copy function, as shown in Fig. 10. This means that if the seedPath is set to a value where the corresponding path does not exist, the exploit fails before reaching the vulnerable sink. This constraint is not easy to capture with symbolic execution, as attempted by Explode.js (Marques et al., 2025), but using execution feedback and an LLM, PoCGen is able to generate a working exploit.
In the first round of exploit generation, PoCGen generates an exploit that calls . After the execution of this exploit, the runtime feedback shows that there was a runtime error happening at the fsExtra.copySync function, which is due to the fact that no directory called "someSeedPath" exists. The LLM then naively tries to fix this by changing "someSeedPath" to "/path/to/valid/seed", which results in the same error. It takes multiple refinements until the LLM requests the definition of shelljs, which allows the reasoning part of the response to understand that it needs to create the directory for a successful exploit. This results in the final exploit shown in Fig. 11.
Figure 11. Example of a successful PoC exploit generation by PoCGen.
async function exploit() {
const initProject = require("n158/src/cli/initProject");
const seedPath = "/path/to/valid/seed";
const targetPath = "targetPath; /usr/bin/genpoc";
const fs = require(’fs’);
if (!fs.existsSync(seedPath)) {
fs.mkdirSync(seedPath, { recursive: true });
const result = await initProject(seedPath, targetPath);
await exploit();
3.6.2. Unsuccessful Example: GHSA-3486-rvxc-hrrj
The vulnerability report GHSA-3486-rvxc-hrrj describes a command injection vulnerability in the gitblame package. The package exports one function, which takes a file path as an argument and after some processing passes it to exec. The relevant parts of the source code are shown in Section 3.6.2.
Figure 12. Example of a command injection vulnerability in the gitblame package.
module.exports = function (file, cb) var dirname = path.dirname(file); var filename = path.basename(file); exec(’git blame ’ + filename, cwd: dirname, ...
PoCGen is not able to generate a working exploit for this vulnerability. The reason is that our validator requires the execution of the command /usr/bin/genpoc. However, the split by path.basename and path.dirname does not allow passing of payloads with "/" to reach the sink without change. This means that our approach is limited by the constraints imposed by the validator. Even with multiple refinements, PoCGen is unable to generate a working exploit, although the exploit for this vulnerability in SecBench.js has the simple payload "& touch gitblame".
4. Threats to Validity
One threat to internal validity is the LLM’s potential to recall exploits from its training data. To mitigate this, we use the same LLM for the AutoGPT baseline, which we would expect to perform similarly well as PoCGen if the results were simply due to memorization. Another threat is that the labeling process of false positive PoC exploits is done manually, which can introduce human bias. To reduce the effects of this bias, we devise a set of objective criteria. For example, prototype pollution cases where the Object.prototype is part of an assignment or a function call are considered false positives. A threat to external validity is that our evaluation is limited to five vulnerability types. However, these vulnerability types correspond to the five most common threat classes in server-side JavaScript (Bhuiyan et al., 2023). Adapting the approach to other vulnerability types would require adapting our prompts, the sinks used by the taint analysis, and the validators. Our approach is implemented for JavaScript and the npm ecosystem, which may limit its applicability to other programming languages and ecosystems. We believe that the core ideas of PoCGen, i.e., how to combine LLMs with program analysis techniques, could be adapted to other contexts, but leave this for future work.
5. Related Work
@@
Greybox fuzzing (Zalewski, 2013; Böhme et al., 2019), applied to source code (Sherman and Nagy, 2025) or binaries (Dinesh et al., 2020), is a common approach for vulnerability detection. To evaluate fuzzing, techniques for reverting fixes (Zhang et al., 2022) and benchmarking methodologies (Klees et al., 2018; Li et al., 2021a) have been proposed. Learning-based vulnerability detection includes neural classification (Li et al., 2018; Fu and Tantithamthavorn, 2022; Chakraborty et al., 2022), graph neural networks (Li et al., 2021b; Steenhoek et al., 2024), and combinations of LLMs with static analysis. Similar to our work, the latter leverages CodeQL to identify taint flows (Li et al., 2024). To support learning-based detection, datasets from commit histories (Zheng et al., 2021) and large-scale vulnerability generation approaches (Nong et al., 2022, 2023) have been introduced. Vulnerability detection is orthogonal to our work, as we assume vulnerabilities are already described in a report but lack an exploit.
@@
The closest work to PoCGen is Explode.js (Marques et al., 2025), which finds vulnerabilities and generates PoC exploits for npm packages. Explode.js uses static dataflow analysis to extract the sequence of function calls required to propagate attacker input to a vulnerable sink. It then applies symbolic execution and constraint solving to generate a PoC exploit. However, Explode.js does not model external functions and libraries during symbolic execution, which limits its effectiveness, as the npm ecosystem heavily relies on small, reusable packages. PoCGen addresses this limitation by leveraging LLMs to generate exploits and by incorporating runtime feedback. LLMs, trained on large code corpora, can better predict the behavior of external functions and reason about inputs that exercise specific program paths.
Other approaches have used symbolic execution to generate exploits for vulnerabilities. FAST (Kang et al., 2023) applies bi-directional dataflow analysis to detect taint paths efficiently, enabling scalable vulnerability detection. It generates exploits by concretizing symbolic path constraints once a vulnerability is found. Node-Medic (Cassel et al., 2023) and Node-Medic-FINE (Cassel et al., 2025) combine dynamic taint analysis with symbolic execution to detect and generate exploits for Node.js packages. Node-Medic-FINE further incorporates fuzzing to generate inputs and explore additional execution paths. All three approaches are outperformed by Explode.js (Marques et al., 2025), which we therefore use as a baseline in our evaluation.
@@
Zhang et al. (Zhang et al., 2023) and Gao et al. (Gao et al., 2025) use LLMs to generate unit tests for Java vulnerabilities given a PoC exploit. Their goal is to encourage developers to update vulnerable dependencies and prevent supply chain attacks. In contrast, our work generates code that directly exploits a vulnerable package, rather than exploiting it through a third-party dependency.
@@
Recent work has explored the potential of LLMs for attacking vulnerable software. PentestGPT (Deng et al., 2024) uses LLMs for penetration testing. Xu et al. (Xu et al., 2024) developed an LLM agent with command-line access to exploit vulnerabilities in Linux and Windows applications. Charan et al. (Charan et al., 2023) investigated using LLMs to generate payloads for exploiting vulnerabilities.
@@
Mitigating vulnerabilities can involve removing unused dependencies (Koishybayev and Kapravelos, 2020) or reducing the privileges of vulnerable code (Vasilakis et al., 2021). Repairing vulnerabilities can be achieved by fine-tuning LLMs to find fixes (Berabi et al., 2024), using LLM agents (Zhang et al., 2024b), or applying generative adversarial networks (GANs) (Harer et al., 2018).
@@
The npm ecosystem faces various security issues, such as injection attacks (Staicu et al., 2018), ReDoS (Staicu and Pradel, 2018), and malicious packages (Sejfia and Schaefer, 2022). Several empirical studies have analyzed npm from a security perspective, including vulnerability propagation (Zimmermann et al., 2019; Liu et al., 2022) and how developers address vulnerabilities (Zahan et al., 2022). Householder et al. (Householder et al., 2020) found that most vulnerability reports lack a public PoC exploit for at least one year. Yadmani et al. (Yadmani et al., 2023) further showed that many PoC exploits on GitHub are themselves malicious. These findings motivate our work on automated PoC exploit generation. To support further research, Bhuiyan et al. (Bhuiyan et al., 2023) introduced the SecBench.js dataset, and Brito et al. (Brito et al., 2023) created the VulcaN dataset.
6. Conclusion
This paper presents PoCGen, an LLM-based approach to automatically generate proof-of-concept exploits for vulnerabilities in npm packages. PoCGen extracts information from the vulnerability report and the codebase to generate a PoC exploit using an LLM. The generated PoC exploits are validated using a set of runtime checkers, and the prompt is refined using static and dynamic information to generate a valid exploit. We evaluated PoCGen on a dataset of 560 vulnerabilities in npm packages. PoCGen generates PoC exploits for 77% of these vulnerabilities, outperforming the state-of-the-art approaches Explode.js and AutoGPT by 35 and 47 absolute percentage points, respectively.
By automating the generation of PoC exploits, PoCGen enables developers and security teams to more rapidly understand and address vulnerabilities, reducing the time between vulnerability disclosure and patch deployment. This automation also improves regression testing, as the generated PoC exploits can be directly used to verify the effectiveness of fixes and to prevent vulnerabilities from reappearing in the future. For security researchers, PoCGen provides an automated way to evaluate the effectiveness of existing mitigation strategies across large sets of vulnerabilities. Furthermore, PoCGen can improve the quality of existing vulnerability reports, including those that are poorly documented or lack existing exploits. Finally, automated PoC generation can facilitate responsible vulnerability disclosure by providing clear, actionable evidence to affected parties, encouraging timely remediation.
@@ 6. Conclusion
The source code of PoCGen, the new dataset, and all experiment scripts are available at https://github.com/sola-st/PoCGen.